# Dissecting Mahalanobis: How Feature Geometry and Normalization Shape OOD Detection

### v1

@@ C.1 Setup and Derivation
Let $z\in\mathbb{R}^{d}\setminus\{0\}$ be a feature vector with the standard Euclidean metric $g_{\mathrm{Euc}}$. We apply a radially symmetric map, $\phi_{\beta}$, to the feature space:
$\phi_{\beta}(z)=\frac{z}{\|z\|^{\beta}},$
where the scalar parameter $\beta\in\mathbb{R}$ controls the degree of radial scaling. This smooth map induces a new Riemannian metric on the domain via the pullback of the Euclidean metric, $g_{\beta}:=\phi_{\beta}^{*}g_{\mathrm{Euc}}$.
The Euclidean metric in polar coordinates $(r,u)$ where $z=ru$, $r=\|z\|$, and $u=z/r\in S^{d-1}$ is $g_{\mathrm{Euc}}=\mathrm{d}r^{2}+r^{2}g_{S^{d-1}}$. Our map transforms the radius to $R=\|\phi_{\beta}(z)\|=r^{1-\beta}$. The differential of the new radius is $\mathrm{d}R=(1-\beta)r^{-\beta}\mathrm{d}r$. By pulling back the Euclidean metric in terms of the new coordinates, we obtain the induced metric $g_{\beta}$:
$g_{\beta}=\mathrm{d}R^{2}+R^{2}g_{S^{d-1}}=\left((1-\beta)r^{-\beta}\mathrm{d}r\right)^{2}+\left(r^{1-\beta}\right)^{2}g_{S^{d-1}}.$
This simplifies to:
$\boxed{g_{\beta}=(1-\beta)^{2}r^{-2\beta}\,\mathrm{d}r^{2}+r^{2(1-\beta)}g_{S^{d-1}},}$
where $g_{S^{d-1}}$ is the standard round metric on the unit sphere. Thus, $\beta$ continuously interpolates between different geometries by re-weighting the radial and angular components of the metric.
C.2 Geometric Interpretation of the Parameter $\beta$
The exponent $\beta$ is a single parameter that continuously interpolates between different geometries by controlling the radial-angular trade-off.
  • Case 1: $\beta=0$ (Euclidean Geometry). The map $\phi_{0}$ is the identity. The induced metric becomes $g_{0}=(1-0)^{2}r^{0}\mathrm{d}r^{2}+r^{2}g_{S^{d-1}}=\mathrm{d}r^{2}+r^{2}g_{S^{d-1}}$, which is precisely the Euclidean metric in polar coordinates.
  • Case 2: $\beta>0$ (Contractive Geometries). This mapping pulls points towards the unit hypersphere. If $0<\beta<1$, radii are compressed but their order is preserved. For $\beta=1$, the radial component of the metric vanishes since $(1-1)^{2}=0$, and the metric reduces to $g_{1}=g_{S^{d-1}}$. This confirms that the space geometrically collapses onto the unit sphere. When $\beta>1$, points that were far from the origin are pulled inward even more strongly, inducing a strong radial contraction. This contractive effect can make the ID distribution more compact.
  • Case 3: $\beta<0$ (Expansive Geometries). Let $\beta=-\gamma$ for $\gamma>0$. The mapping becomes $\phi_{-\gamma}(z)=z\cdot\|z\|^{\gamma}$. This transformation pushes points with large norms even further from the origin, radially stretching the space. This may be optimal when ID class manifolds are already well-separated and situated far from the origin, as it further increases inter-cluster distances while creating a large void around the origin where OOD samples can be easily detected.
C.3 Mahalanobis Distance with Radially Scaled $\ell_{2}$ Normalization
The final OOD detector combines this geometric transformation with the standard Mahalanobis distance calculation. The procedure is as follows: first, a whitening transformation is applied to the features. Then, we apply the radially scaled $\ell_{2}$ normalization map $\phi_{\beta}$ to the whitened features. Finally, we compute the standard Mahalanobis distance (which, in the whitened space, is equivalent to the squared Euclidean distance) between the mapped test feature and the mapped class mean. The final OOD score is the minimum distance to any class mean:
$\text{Score}(z)=\min_{c}D_{M}(\phi_{\beta}(z^{\prime}),\phi_{\beta}(\mu^{\prime}_{c})),$
where $z^{\prime}$ and $\mu^{\prime}_{c}$ are the whitened versions of the feature and class mean, and $D_{M}$ is the Mahalanobis distance. By tuning $\beta$, we directly control the geometry of the space where the distance is measured, providing a principled way to align the feature distribution with the assumptions underlying Mahalanobis-based OOD detection.
Appendix D Full results
D.1 Cross-model performance
@@
D.2 Detailed OOD performance
The results demonstrate that conformal normalization significantly improves OOD detection performance across diverse vision model architectures. Among the tested methods, the optimal variants of Mahalanobis Distance (MD*) and Relative Mahalanobis Distance (RMD*) consistently achieve the lowest false positive rates, with RMD* performing particularly well on transformer-based models like ViT and DeiT3. The Mahalanobis-based variants are defined as follows: MD* uses the empirically optimal $\beta$; $\hat{\textbf{MD}}$ uses the regression-predicted $\hat{\beta}$; MD (standard) fixes $\beta=0$; and MD++ (Mahalanobis++) fixes $\beta=1$. The performance gains are most pronounced on larger models (e.g., ViT-L, DeiT3-L) and those pre-trained on larger datasets (In21k, In22k), suggesting that conformal normalization is especially beneficial for models with richer feature representations. Traditional methods like MSP and MLS show relatively poor performance, while the proposed conformal variants provide substantial improvements in OOD detection reliability.
Table 3: False positive rate (FPR, $\downarrow$) and validation accuracy (Acc) across different models and detectors. In most cases, the optimal MD* or RMD* achieves the best overall performance. Note that higher accuracy does not necessarily correspond to a lower FPR.
Model Acc MSP MLS KNN VIM MD* $\hat{\text{MD}}$ MD MD++ RMD* $\hat{\text{RMD}}$ RMD RMD++ BEiTV2 FT In1k 85.5 0.522 0.507 0.426 0.393 0.365 0.375 0.402 0.376 0.363 0.367 0.391 0.373 BEiTV2 FT In21k 85.1 0.380 0.258 0.351 0.295 0.244 0.274 0.436 0.298 0.312 0.324 0.331 0.325 MAE FT In1k 83.5 0.542 0.557 0.444 0.406 0.392 0.399 0.435 0.403 0.382 0.394 0.417 0.393 DINOV2 83.0 0.449 0.328 0.402 0.289 0.326 0.334 0.338 0.345 0.410 0.412 0.414 0.412 ViT 77.1 0.565 0.504 0.500 0.530 0.445 0.456 0.457 0.454 0.442 0.443 0.449 0.446 ViT In21K 75.6 0.511 0.373 0.642 0.340 0.493 0.518 0.532 0.577 0.378 0.392 0.398 0.398 ViT-S In21K In1k 75.8 0.572 0.444 0.538 0.412 0.398 0.407 0.510 0.416 0.442 0.448 0.449 0.452 ViT In21K In1k 78.5 0.537 0.407 0.477 0.361 0.327 0.346 0.357 0.387 0.365 0.375 0.376 0.376 ViT-L In21K In1k 83.6 0.448 0.298 0.343 0.250 0.229 0.239 0.253 0.282 0.259 0.268 0.269 0.269 DeiT3 83.5 0.550 0.592 0.475 0.472 0.425 0.428 0.433 0.430 0.386 0.390 0.408 0.399 DeiT3 FB In22k In1k 83.8 0.609 0.649 0.411 0.398 0.381 0.388 0.407 0.388 0.389 0.395 0.405 0.392 DeiT3 In21k In1k 85.0 0.567 0.643 0.370 0.375 0.348 0.356 0.376 0.356 0.331 0.338 0.366 0.351 DeiT3-L In22k In1k 85.7 0.581 0.659 0.359 0.397 0.320 0.333 0.366 0.342 0.335 0.347 0.376 0.359 EVA02 82.0 0.491 0.395 0.547 0.388 0.378 0.401 0.534 0.446 0.426 0.442 0.440 0.440 EVA02 FT In1k 84.2 0.532 0.553 0.406 0.439 0.368 0.374 0.376 0.374 0.392 0.393 0.403 0.398 EVA02 FT In21k 80.2 0.446 0.343 0.550 0.329 0.468 0.498 0.565 0.506 0.358 0.368 0.367 0.377 EVA02-S FT In22k In1k 82.2 0.592 0.648 0.442 0.448 0.404 0.409 0.422 0.410 0.425 0.426 0.443 0.436 EVA02 FT In21k In1k 82.2 0.530 0.589 0.423 0.371 0.369 0.383 0.408 0.382 0.373 0.383 0.403 0.391 EVA02-L FT In22k In1k 84.8 0.438 0.430 0.384 0.369 0.351 0.368 0.370 0.378 0.308 0.314 0.326 0.318 ViT CLIP In1k 84.7 0.552 0.653 0.411 0.417 0.364 0.386 0.402 0.382 0.366 0.370 0.403 0.386 ViT CLIP In12k In1k 85.4 0.490 0.519 0.325 0.300 0.246 0.254 0.335 0.278 0.294 0.297 0.325 0.309 ViT-L CLIP In12k In1k 86.1 0.450 0.436 0.301 0.280 0.258 0.278 0.297 0.271 0.259 0.264 0.293 0.277

### v3

@@ D.1 Setup and induced metric
Let $z\in\mathbb{R}^{d}\setminus\{0\}$ be a feature vector with Euclidean metric $g_{\mathrm{Euc}}$. Consider the radial map
$\phi_{\beta}(z)\;=\;\frac{z}{\|z\|^{\beta}},\qquad\beta\in\mathbb{R},$ (30)
which preserves direction but rescales radius. Writing z=ruz=r\,u with $r=\|z\|$ and $u\in S^{d-1}$, the Euclidean metric decomposes as
$g_{\mathrm{Euc}}\;=\;\mathrm{d}r^{2}+r^{2}g_{S^{d-1}}.$ (31)
Under $\phi_{\beta}$, the radius becomes $R=r^{1-\beta}$ and $\mathrm{d}R=(1-\beta)\,r^{-\beta}\mathrm{d}r$. The pullback metric $g_{\beta}\triangleq\phi_{\beta}^{\ast}g_{\mathrm{Euc}}$ therefore satisfies
$g_{\beta}\;=\;(1-\beta)^{2}r^{-2\beta}\,\mathrm{d}r^{2}\;+\;r^{2(1-\beta)}g_{S^{d-1}}.$ (32)
Eq. (32) makes explicit that $\beta$ changes the relative weighting of radial and angular variations. This is the geometric reason that empirical quantities computed after $\phi_{\beta}$ (e.g., covariance spectra and neighborhood distances used for LID) vary smoothly with $\beta$.
D.2 How this interacts with Mahalanobis-style scoring
In the main paper we apply $\phi_{\beta}$ to features and then fit the (tied) Gaussian statistics used by Mahalanobis variants. Equivalently, $\phi_{\beta}$ changes the distribution of deviations $\delta_{\beta}(z)$ before inserting them into a quadratic score $S_{\beta}(z)=\delta_{\beta}(z)^{\top}\Sigma(\beta)^{-1}\delta_{\beta}(z)$. Thus, $\beta$ does not define a new detector; it defines a one-parameter family of geometrically deformed representations. Through Eq. (32), this deformation changes: (i) neighborhood structure (hence $m(\beta)$), (ii) scatter spectra (hence $s(\beta)$), and (iii) how quadratic weighting amplifies directional deviations.
D.3 Interpreting $\beta$
Eq. (32) yields a compact interpretation of typical regimes:
  • $\beta=0$ (identity). $\phi_{\beta}$ is the identity and $g_{\beta}=g_{\mathrm{Euc}}$.
  • $0<\beta<1$ (moderate contraction). Radii shrink as $r^{1-\beta}$ while angular structure is retained.
  • $\beta=1$ (spherical projection). Radii become constant ($R=1$): radial variability is removed and only angles remain.
  • $\beta>1$ (strong contraction). Large radii are aggressively compressed; this can suppress variability but may also collapse useful structure.
  • $\beta<0$ (expansion). Radii are amplified; norm differences become more prominent, which can increase sensitivity to radial outliers.
D.4 $\beta$ selection from the proxy curve
For each model we evaluate the proxy
$P(\beta)=m(\beta)\,|s(\beta)|$ (33)
on a discrete grid $\mathcal{B}=\{\beta_{1}<\cdots<\beta_{T}\}$. Empirically, $P(\beta)$ is usually inverted-U shaped (so maximizing $P$ is the typical behavior), but some representations can produce U-shaped curves (where minimizing $P$ is appropriate). Rather than hard-coding “maximize” versus “minimize,” we select the most pronounced interior turning point.
Interior turning point rule.
Let $P_{t}\triangleq P(\beta_{t})$. We first define a simple endpoint reference level
$P_{\mathrm{end}}\;\triangleq\;\frac{P_{1}+P_{T}}{2}.$
Among interior grid points $t\in\{2,\ldots,T-1\}$, we choose the one farthest from this endpoint level:
$\hat{t}\in\arg\max_{t\in\{2,\ldots,T-1\}}\big|P_{t}-P_{\mathrm{end}}\big|,\qquad\hat{\beta}\triangleq\beta_{\hat{t}}$. (34)
This rule returns the interior maximum when the curve is inverted-U, and the interior minimum when the curve is U-shaped, while discouraging endpoint-driven choices.
Nearly monotone curves.
If the proxy has no clear interior turning point on the finite grid (i.e., the largest interior deviation is very small), we fall back to the default behavior and set
$\hat{\beta}\in\arg\max_{\beta\in\mathcal{B}}P(\beta).$
In our experiments, this situation is rare and typically occurs when the true extremum lies outside the evaluated range.
Connection to the main text.
The selection rule is intentionally agnostic to whether $P(\beta)$ should be maximized or minimized. In Sec. 7.1 we further explain why both U-shaped and inverted-U regimes occur by relating the proxy curve to the instability functionals $\mathcal{I}(\beta)$ and its low-dimensional approximation $\widehat{\mathcal{I}}(\beta)$.
Appendix E Unified Stability Theory: Definitions, Mechanism, and Empirical Evidence
This appendix provides the formal definitions used by Unified Stability Theory (UST), derives the stretch mechanism underlying the main text, and summarizes the supporting empirical evidence. Unless stated otherwise, all quantities are computed on last-layer features and estimated using ID data only.
E.1 Quadratic scores and the instability functional
Each Mahalanobis-style detector specifies a centered deviation $\delta_{\beta}(z)$, where $z$ is a feature vector, and a tied scatter (or covariance) estimate $\Sigma(\beta)$ computed from ID data. We assume $\Sigma(\beta)\succ 0$ for all $\beta$ under consideration, ensuring $\Sigma(\beta)^{-1}$ exists, and that $\|\delta_{\beta}(z)\|>0$ almost surely under $\mathcal{D}_{\mathrm{ID}}$ so $\log\|\delta_{\beta}(z)\|^{2}$ is well-defined.
The corresponding quadratic score is
$S_{\beta}(z)\;\triangleq\;\delta_{\beta}(z)^{\top}\Sigma(\beta)^{-1}\delta_{\beta}(z).$ (35)
We quantify ID-side sensitivity of the score using
$\mathcal{I}(\beta)\;\triangleq\;\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log S_{\beta}(z)\big].$ (36)
Exact factorization and channel decomposition.
For $\delta_{\beta}(z)\neq 0$, define
$W_{\beta}(z)\;\triangleq\;\frac{\delta_{\beta}(z)^{\top}\Sigma(\beta)^{-1}\delta_{\beta}(z)}{\|\delta_{\beta}(z)\|^{2}}.$ (37)
Then
$S_{\beta}(z)=\|\delta_{\beta}(z)\|^{2}\,W_{\beta}(z),\qquad\log S_{\beta}(z)=\log\|\delta_{\beta}(z)\|^{2}+\log W_{\beta}(z).$ (38)
We verify the factorization in Eq. (38) numerically to machine precision (Table 2). Taking the variance over ID samples yields the exact decomposition
$\mathcal{I}(\beta)=A_{\delta}(\beta)+A_{W}(\beta)+2A_{\times}(\beta),$ (39)
where
$A_{\delta}(\beta)$ $\triangleq\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log\|\delta_{\beta}(z)\|^{2}\big],$ (40)
$A_{W}(\beta)$ $\triangleq\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log W_{\beta}(z)\big],$
$A_{\times}(\beta)$ $\triangleq\operatorname{Cov}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big(\log\|\delta_{\beta}(z)\|^{2},\log W_{\beta}(z)\big).$
In the main text we refer to $A_{\delta}$ as the size channel and $A_{W}$ as the stretch channel. Across models and $\beta$, both channels $A_{\delta}(\beta)$ and $A_{W}(\beta)$ exhibit nontrivial variation, and the cross-term $A_{\times}(\beta)$ is predominantly negative (Fig. 11), indicating partial compensation between size and stretch.
Empirical evidence.
We numerically verify Eq. (38) to machine precision (Table 2), and we observe substantial variation in both $A_{\delta}(\beta)$ and $A_{W}(\beta)$ across $\beta$ and models (Figure 11). Moreover, the cross-term is systematically negative in our configurations, indicating partial compensation between size and stretch (also summarized in Figure 11).
Table 2: Numerical residual for the factorization $S_{\beta}(z)=\|\delta_{\beta}(z)\|^{2}W_{\beta}(z)$ over all evaluated configurations.
Statistic max rel. err mean rel. err #configs
All 7.34e-07 4.56e-08 462
Figure 11: Distribution of the three terms $(A_{\delta},A_{W},A_{\times})$ across models and $\beta$. The cross-term is predominantly negative, consistent with compensation between size and stretch.
E.2 Stretch mechanism: spectrum $\times$ allocation geometry
This subsection derives the representation of $W_{\beta}$ that motivates the intervention experiments.
Write the eigendecomposition $\Sigma(\beta)=U(\beta)\Lambda(\beta)U(\beta)^{\top}$, where $\Lambda(\beta)=\mathrm{diag}(\lambda_{1}(\beta),\ldots,\lambda_{d}(\beta))$ with $\lambda_{1}(\beta)\geq\cdots\geq\lambda_{d}(\beta)>0$, and $U(\beta)=[u_{1}(\beta),\ldots,u_{d}(\beta)]$ is an orthonormal eigenbasis.
Define the allocation of a deviation $\delta_{\beta}(z)$ onto eigen-directions by
$p_{i}(z;\beta)\triangleq\frac{(u_{i}(\beta)^{\top}\delta_{\beta}(z))^{2}}{\|\delta_{\beta}(z)\|^{2}},\qquad\sum_{i=1}^{d}p_{i}(z;\beta)=1.$ (41)
When $\beta$ is clear from context we abbreviate $p_{i}(z;\beta)$ as $p_{i}(z)$. Substituting the eigendecomposition into Eq. (37) gives the exact identity
$W_{\beta}(z)=\sum_{i=1}^{d}\frac{p_{i}(z)}{\lambda_{i}(\beta)}.$ (42)
Interpretation.
Equation (42) shows that pointwise exposure is controlled jointly by (i) spectral heterogeneity through $\{\lambda_{i}(\beta)\}$ and (ii) allocation geometry through the random simplex vector $p(z)$. In particular, variability in $W_{\beta}$ (and hence $A_{W}$) can be reduced either by flattening the spectrum (e.g., shrinkage, spectral shaping) or by making allocations more uniform across samples (e.g., reducing dominance/spikiness or explicitly smoothing $p(z)$).
E.3 Interventions isolating size and stretch
We now summarize the intervention experiments that isolate the two channels in Eq. (39). Each intervention targets a single mechanism suggested by Eq. (42).
Rotation changes allocation geometry while preserving norms.
Applying a random orthogonal map to deviations, $\delta_{\beta}(z)\mapsto Q\,\delta_{\beta}(z)$ with $Q^{\top}Q=I$, leaves $\|\delta_{\beta}(z)\|$ unchanged (up to numerical precision) but alters the projections $\{u_{i}(\beta)^{\top}\delta_{\beta}(z)\}_{i=1}^{d}$. Hence, the allocation vector $p(z)$ in Eq. (42) is redistributed across eigen-directions, which directly perturbs the exposure term $W_{\beta}(z)$. Empirically, the rotation intervention induces large changes in $A_{W}$ while $\Delta A_{\delta}$ stays near zero across configurations (Fig. 12).
Figure 12: Rotation intervention: paired changes $(\Delta A_{\delta},\Delta A_{W})$ under orthogonal mixing. $\Delta A_{\delta}$ concentrates near zero (numerical scale $\sim 10^{-7}$), whereas $\Delta A_{W}$ varies substantially, indicating that rotation primarily perturbs stretch variability.
Allocation smoothing collapses stretch variability.
We construct a controlled family of allocations
$p^{(\eta)}(z)=(1-\eta)p(z)+\eta\,\bar{p},\qquad\eta\in[0,1],$ (43)
where $\bar{p}$ is the ID-average allocation. Substituting $p^{(\eta)}$ into Eq. (42) removes point-dependent exposure as $\eta\to 1$. Empirically, $A_{W}(\eta)$ decreases monotonically and approaches numerical zero at $\eta=1$ (Figure 13).
Figure 13: Allocation smoothing: $A_{W}(\eta)$ decreases and collapses as $\eta\to 1$, consistent with Eq. (42).
Spectrum shaping changes $A_{W}$ with allocations held fixed.
We modify the eigenvalues while keeping the eigenvectors fixed (e.g., $\lambda_{i}^{\prime}\propto\lambda_{i}^{\alpha}$ with trace renormalization) and recompute $W_{\beta}$ using the same allocations $p(z)$ in Eq. (42). This directly tests the spectral factor in the exposure mechanism and yields large changes in $A_{W}$.
Radial rescaling leaves $W$ invariant under fixed $\Sigma$.
If $\Sigma(\beta)$ is fixed and deviations are rescaled as $\delta^{\prime}(z)=c(z)\delta(z)$ for any $c(z)>0$, then
$W^{\prime}(z)=\frac{\delta^{\prime}(z)^{\top}\Sigma(\beta)^{-1}\delta^{\prime}(z)}{\|\delta^{\prime}(z)\|^{2}}=\frac{c(z)^{2}\,\delta(z)^{\top}\Sigma(\beta)^{-1}\delta(z)}{c(z)^{2}\,\|\delta(z)\|^{2}}=W(z).$
Therefore $A_{W}=\operatorname{Var}(\log W)$ is invariant under such rescalings when $\Sigma(\beta)$ is held fixed. We verify this fixed-$\Sigma$ invariance numerically in Table 3. For completeness, we also report a refit-$\Sigma$ variant as a sensitivity analysis, where invariance is not expected. We include this invariance as a sanity check to ensure that our size-only manipulations cannot affect the exposure term $W$ unless the whitening operator $\Sigma^{-1}$ itself changes.
Table 3: Scale invariance of $A_{W}$ under fixed $\Sigma$.
Metric mean $|\Delta A_{W}|$ #configs
E8 $10^{-6}$ 1386
E.4 Low-dimensional proxy and the role of $m_{k}$
This subsection supports the claim in Sec. 7.4 that, within a fixed representation, the ID instability trajectory $\beta\mapsto\mathcal{I}(\beta)$ can be captured by a small set of ID-only geometry statistics. We consider the additive proxy
$\widehat{\mathcal{I}}(\beta)\;=\;a\,\log m_{k}(\beta)\;+\;b\,|s(\beta)|\;+\;c\,\mathrm{curv}(\beta),$ (44)
where $m_{k}(\beta)$ is the kkNN LID estimate in $\phi_{\beta}$-space and $(|s(\beta)|,\mathrm{curv}(\beta))$ summarize the eigenspectrum of the relevant scatter matrix (e.g., $\Sigma(\beta)$ or $S_{w}(\beta)$).
Spectral descriptors.
Let $\lambda_{1}(\beta)\geq\cdots\geq\lambda_{d}(\beta)>0$ denote the ordered eigenvalues and define the log-spectrum $\ell_{i}(\beta)\triangleq\log\lambda_{i}(\beta)$. We fit an affine model
$\ell_{i}(\beta)\;\approx\;\alpha(\beta)+s(\beta)\,i,\qquad i=1,\ldots,d,.$
by least squares and report:
  • Slope. $s(\beta)$ is the fitted slope; we use $|s(\beta)|$ as a scale-free measure of overall spectral decay (steeper decay $\Rightarrow$ larger $|s|$).
  • Curvature (residual curvature). We measure nonlinearity of the log-spectrum by the lack-of-fit of the same linear regression:
    $\mathrm{curv}(\beta)\;\triangleq\;1-R^{2}(\beta),$
    so $\mathrm{curv}(\beta)\approx 0$ when $\log\lambda_{i}(\beta)$ is close to affine in $i$ (near-exponential eigen-decay), and larger values indicate a more non-affine/irregular spectrum. Intuitively, this is the fraction of log-spectrum variance not explained by a straight line.
Per-model least squares (ID-only).
Fix a representation (model) and evaluate the proxy components on a discrete grid $\mathcal{B}=\{\beta_{1},\ldots,\beta_{T}\}$ using ID features only. Define
$x(\beta_{t})=\big[\log m_{k}(\beta_{t}),\ |s(\beta_{t})|,\ \mathrm{curv}(\beta_{t})\big]\in\mathbb{R}^{3},\qquad y(\beta_{t})=\mathcal{I}(\beta_{t}).$
We fit $(a,b,c)$ by least squares without an intercept:
[…]
We omit an intercept since we evaluate agreement across $\beta$ (e.g., via rank correlation); an additive constant would not change the ordering of $\widehat{\mathcal{I}}(\beta)$, and the proxy is intended to capture the shape of the instability trajectory.
Proxy tracking along the $\beta$ trajectory.
Within each model, $\widehat{\mathcal{I}}(\beta)$ tracks $\mathcal{I}(\beta)$ over the $\beta$ sweep with high rank agreement (Table 4; main-paper visualization: the overlay plot in Figure 9).
Coupling beyond spectrum.
Regressing $A_{W}(\beta)$ on $(|s(\beta)|,\mathrm{curv}(\beta))$ and then adding $\log m_{k}(\beta)$ yields a consistent improvement in explained variance across models (Median $\Delta R^{2}\geq 0.05$; mean $\Delta R^{2}$ is typically higher, with a wide IQR, indicating heterogeneous gains across model families). This supports treating $m_{k}$ as an additional ID geometry statistic that modulates exposure-driven stretch variability.
Table 4: Proxy tracking: within-model rank correlation between $\widehat{\mathcal{I}}(\beta)$ and $\mathcal{I}(\beta)$ over $\beta$.
$k$ median Spearman $\rho$ IQR
10 0.532 0.670
25 0.536 0.670
50 0.540 0.670
100 0.547 0.663
E.5 RMD: Term-wise Unified Stability Lens
RMD is commonly implemented as a difference of two Mahalanobis-type quadratic forms, $S_{\mathrm{RMD},\beta}(z)=S^{(1)}_{\beta}(z)-S^{(2)}_{\beta}(z)$, where the deviations $\delta^{(j)}_{\beta}$ and scatter estimates $\Sigma^{(j)}(\beta)$ may differ between terms. As a result, RMD does not generally admit the single-$\Sigma$ representation in Eq. (7).
E.6 Exact term-wise factorization and channels
For each term $j\in\{1,2\}$, define
$S^{(j)}_{\beta}(z)=\delta^{(j)}_{\beta}(z)^{\top}\Sigma^{(j)}(\beta)^{-1}\delta^{(j)}_{\beta}(z),\qquad W^{(j)}_{\beta}(z)=\frac{S^{(j)}_{\beta}(z)}{\|\delta^{(j)}_{\beta}(z)\|^{2}}.$
Then the exact factorization holds term-wise:
$S^{(j)}_{\beta}(z)=\|\delta^{(j)}_{\beta}(z)\|^{2}\,W^{(j)}_{\beta}(z),\qquad\log S^{(j)}_{\beta}(z)=\log\|\delta^{(j)}_{\beta}(z)\|^{2}+\log W^{(j)}_{\beta}(z),$
and the instability functional for each term decomposes as
$\mathcal{I}^{(j)}(\beta)\triangleq\operatorname{Var}[\log S^{(j)}_{\beta}]=A^{(j)}_{\delta}(\beta)+A^{(j)}_{W}(\beta)+2A^{(j)}_{\times}(\beta),$
with the obvious definitions of $A^{(j)}_{\delta},A^{(j)}_{W},A^{(j)}_{\times}$.
Appendix F Full results
F.1 Cross-model performance
@@
F.2 Detailed OOD performance
Table 6 reports FPR@95 ($\downarrow$) for each (model, detector) pair together with the corresponding validation accuracy. Across a wide range of backbones, the radially scaled (RS) variants provide the most reliable gains within the Mahalanobis families: whenever RS-MD or RS-RMD is compared against its direct baselines (MD++/MD or RMD++/RMD), it frequently attains a lower FPR@95, indicating that radially scaled normalization is a strong and generally beneficial modification. Although a few models still achieve their absolute best score with non-Mahalanobis detectors (e.g., MLS or VIM in isolated cases), the RS variants remain consistently competitive and, in aggregate, deliver the lowest mean FPR@95 across models. This pattern suggests that the RS mechanism improves robustness across heterogeneous pretrained representations, whereas accuracy alone is not predictive of OOD performance; high-accuracy models can still exhibit higher FPR@95 than lower-accuracy ones, motivating direct evaluation of OOD detection rather than using accuracy as a proxy.
Table 6: FPR@95 ($\downarrow$) and validation accuracy (Acc, %) across models and detectors. Bold denotes the lowest FPR@95 within each model across all detectors. Light green highlights cases where RS-MD outperforms both MD++ and MD, and where RS-RMD outperforms both RMD++ and RMD; in these cases, the corresponding baselines are grayed out. The Avg row reports the mean FPR@95 across models for each detector.
[…]